### v2

# Unraveling Token Prediction Refinement and Identifying Essential Layers in Language Models

Abstract
This research aims to unravel how large language models (LLMs) iteratively refine token predictions through internal processing. We utilized a logit lens technique to analyze the model’s token predictions derived from intermediate representations. Specifically, we focused on (1) how LLMs access and utilize information from input contexts, and (2) how positioning of relevant information affects the model’s token prediction refinement process. On a multi-document question answering task with varying input context lengths, we found that the depth of prediction refinement (defined as the number of intermediate layers an LLM uses to transition from an initial correct token prediction to its final, stable correct output), as a function of the position of relevant information, exhibits an approximately inverted U-shaped curve. We also found that the gap between these two layers, on average, diminishes when relevant information is positioned at the beginning or end of the input context. This suggested that the model requires more refinements when processing longer contexts with relevant information situated in the middle. Furthermore, our findings indicate that not all layers are equally essential for determining final correct outputs. Our analysis provides insights into how token predictions are distributed across different conditions, and establishes important connections to existing hypotheses and previous findings in AI safety research and development.
1 Introduction
Recent advances in technologies have seen an increasing number of applications of capable AI systems in various domains. One example is that of large language models (LLMs) that have exhibited improved capabilities in content and code generation and language translation. As LLMs rapidly advance in sophistication and generality, understanding them is essential to ensure their alignment with human values and prevent catastrophic outcomes (Hendrycks and Mazeika, 2022; Hendrycks et al., 2023; Ji et al., 2024). Research that aims to improve an understanding of LLM goes beyond simple performance metrics, to unravel the internal workings of LLMs. This work utilized an observational approach that analyzes the inner workings of neural networks: Logit Lens (as discussed in detail in Section 2). We focused on how LLMs access and use information from input contexts, and how positioning of relevant information affects the model’s token prediction refinement process, to gain more insights that potentially aid the development of methods that can ensure trustworthiness in LLMs.
Mechanistic interpretability (MI) is a growing approach that aims to precisely define computations within neural networks. It seeks to fully specify a neural network’s computations, aiming for a granular understanding of model behavior, akin to reverse-engineering the model’s processes into pseudo-code. This approach stands out for its ambitious goal of complete reverse engineering and its focus on AI safety. (for a list of introduction and comprehensive review, see (Olah, 2022; Nanda, 2022; Olah et al., 2020; Sharkey et al., 2022; Olah et al., 2018; Nanda, 2023, 2024).)
MI uses a variety of tools, from observational analysis to causal interventions (for a more detailed discussion of taxonomies of explainable artificial intelligence and MI methods, see (Das et al., 2020; Speith, 2022; Bereska and Gavves, 2024; Zytek et al., 2022; Zhang et al., 2021)). Causal methods span from purely observational, which analyze existing representations without manipulation, to interventional approaches that directly perturb model components with the hope to reveal causal relationships. In terms of learning phases, post-hoc techniques are applied after training, while intrinsic methods are designed to enhance interpretability during the training process itself. This is to gain a deeper understanding of the behavior of the model from the beginning.
Another approach, given the scope of analysis and application, is to categorize methods based on general tendencies of interpretability. Some can provide both local and global insight, or partial and comprehensive understanding. Probing techniques (Alain, 2016; Ettinger et al., 2016; Hupkes et al., 2018), for instance, range from local to global: simple linear probes reveal individual feature insights, while advanced structured probes (Burns et al., 2022) expose broader patterns. Sparse autoencoders (Cunningham et al., 2023), while decomposing individual neuron activations (local), aim to disentangle features throughout the model (global). Path patching (Wang et al., 2022; Goldowsky-Dill et al., 2023) transforms local interventions into global understanding by tracing information flow across layers, showing how small perturbations can illuminate the entire model’s behavior.
In this work, logit lens technique–one among the other observational methods–was employed (1) to examine how network layers iteratively refine token predictions, and (2) to observe its relationships to input context lengths and the position of relevant information within input context, and to another method under the same category: probing. In the remainder of this paper, details about logit lens technique is provided in Section 2. Section 3 presents the relationships between logit lens and input context lengths and the position of relevant information within the input context. Relation of logit lens to probing is discussed in Section 4. Section 5 gives an overview of how understanding inner working mechanisms of LLM is relevant to AI safety research and development.
2 Logit Lens: A View into Next Token Information
Logit lens technique, as first introduced by nostalgebraist (2020), provides a way for examining intermediate representations of model’s predictions at different stages of the model’s output generation process, thereby providing insights into how the model’s internal representations evolve as model layers iteratively refine predictions.
Specifically, given a pre-trained large language model (LLM), we use logit lens to decode a probability distribution over tokens from each intermediate layer. These token distributions represent model predictions after $l\in\{1,..,L\}$ layers of input processing. Given a sequence of tokens $t_{1},..,t_{n}\in V$, and $h_{l}^{(i)}\in R^{d}$ denoting the hidden state of token $t_{i}$ at layer $l$, the logits of the predictive distribution $p(t_{n+1}|t_{1},..,t_{n})$ are given by
$[\text{logit}_{1},...,\text{logit}_{|V|}]=W_{U}\cdot\text{LayerNorm}_{L}(h_{L}^{(n)}),$ (1)
where $W_{U}\in R^{|V|\times d}$ denotes an embedding matrix, and $\text{LayerNorm}_{L}$ is the pre-embedding layer normalization. The logit lens applied the same unembedding operation to the earlier hidden states $h_{l}^{(i)}$:
$[\text{logit}_{1}^{l},...,\text{logit}_{|V|}^{l}]=W_{U}\cdot\text{LayerNorm}_{L}(h_{l}^{(n)}),$ (2)
where an intermediate predictive distribution over tokens at layer $l$, $p_{l}(t_{n+1}|t_{1},..,t_{n})$, can be obtained.
In this experiment, we utilized logit lens to hidden states of GPT-2 processing data of internet. For example, given an instance "Hinton is a prominent figure in the field of artificial intelligence and deep learning.", shown in Figure 1, we observed how distributions over next token predictions gradually converge to the final distribution. The observed, general trend is that initial predictions of tokens resemble plausible completions but seems far from the correct prediction. As layers progress, the predictions become more accurate with respect to the ground truth (next token). Roughly in the middle layers, a model forms a correct token of the next token, with later layers refining these predictions with higher predicted probabilities.
(a) Maximum Probability
(b) Cross-Entropy
(c) Forward Kullback-Leibler (KL) Divergence
Figure 1: The logit lens were applied to the hidden states of GPT-2 processing Hinton is a prominent figure in the field of artificial intelligence and deep learning. The vertical axis represents layers ranging from 0 - 12 (bottom to top), and the horizon axis represents input tokens. Each cell illustrates the most likely next tokens (i.e., top-1 token prediction) in a sequence that the respective hidden state predicts. In case of maximum probability, the darker cells correspond to higher probabilities. And in case of cross-entropy and forward Kullback–Leibler (KL) divergence, the lighter cells correspond to higher probabilities.
Take the 6th and 9th columns of Figure 1 as an example of the case where linking words (e.g., conjunctions and subordinating conjunctions), occur, in this case, "figure in" and "field of". The model predicts the next word correctly in the early layer, without changing to other tokens in the later layers. And the model’s confidence evolves over layers to the degree where the model is more confident about its final token prediction (nearly 100% of predicted probabilities). Despite being trained on vast amounts of text without explicit syntactic rules, language models, GPT-2 in this case, show an ability to capture dependency relations, and structural aspects of language. Study reported by (Hewitt and Manning, 2019) shows that language models capture hierarchical syntactic structures, such as nested clauses, phrase boundaries, and subject-verb-object relationships. However, an intriguing question is how transformer-based models, trained solely on natural language data, manage to learn its hierarchical structure and generalize to sentences with unseen syntactic patterns?–despite lacking any explicitly encoded structural bias.
(Ahuja et al., 2024) explored the inductive biases in transformers that may drive this generalization. Through their extensive experiments on synthetic datasets with various training objectives, they found that, while objectives like sequence-to-sequence and prefix language modeling often fail to produce hierarchical generalization, models trained with the language modeling objective consistently are able to produce hierarchical generalization. The authors also conducted experiments on model pruning to examine how transformers encode hierarchical structure. They found (pruned) sub-networks with distinct generalization behaviors aligned with hierarchical structures and with linear order. Model pruning involves reducing the size of networks by removing unnecessary parameters while maintaining performance, Pruning models not only reduces model complexity and computational costs but also provides insights into model efficiency and interpretability. For example, pruning can reveal which parts of a network are most critical for certain tasks, shedding light on how models work internally.
Another interesting aspect is the model’s demonstrated ability to grasp compositional understanding. For instance, in columns 11st and 14th, it recognizes how meaning arises from combining parts of a sentence, such as "artificial … intelligence" and "deep … learning," despite having multiple word choices that could follow "artificial" or "deep." Given the vast token space available at each layer, it is intriguing that the model not only selects the correct token but also reflects uncertainty in its decision with low probabilities. This somehow highlights its nuanced handling of language structure and meaning. However, this ability may not extend to more complex cases where the length of the input context and the position of relevant information vary. In such scenarios, as shown in Section 3 for preliminary results, the model’s performance may struggle to maintain the same level of compositional understanding and accuracy.
Figure 2: Left: Shifting the location of relevant information within the model’s input context reveals an U-shaped performance pattern. Model performs best when the relevant information is at the very beginning or end of the input, while its performance on average declines significantly when the critical information appears in the middle of the context. Right: Similarly, changing the position of relevant information produces a nearly inverted-U pattern, illustrating when the top-1 token first aligns with the correct answer and when it is finalized. The gap between these two layers on average is smaller when the relevant information is placed at the beginning or end of the input, indicating that the model needs more prediction refinements when processing long input context with the relevant information positioned in the middle. The experiments were repeated for 10 independent runs, and dots represent mean (arithmetic average) performance, with shaded bands represent 95% confidence interval (CI).
3 Relation of Logit Lens to Input Context Lengths and the Position of Relevant Information within Input Context
Input contexts for LLMs can contain thousands of tokens, especially when processing lengthy documents or integrating external information. For these tasks, LLMs must efficiently handle long sequences. Study done by (Ivgi et al., 2023; Liu et al., 2024) explored how LLMs use input contexts to perform downstream tasks.
(Liu et al., 2024) designed an experiment to assess how LLMs access and use information from input contexts. They manipulated two factors: (1) the length of the input context and (2) the position of relevant information within it, to evaluate their impact on model performance. They hypothesized that if LLMs can reliably use information from long contexts, performance should remain stable regardless of the position of relevant information.
Multi-document question answering was selected for the experiments, where models must reason over multiple documents to extract relevant information and answer a question. This task reflects the retrieval-augmented generation process used in various applications. In more details about the set up, they controlled two key factors: (i) input context length by varying the number of documents (simulating different retrieval volumes), and (ii) the position of relevant information by altering the document order, placing the relevant one at the beginning, middle, or end of the context. Their findings showed that the position of relevant information in the input context significantly impacts model performance, revealing that current language models struggled to consistently access and use information in long contexts. Notably, they observed a U-shaped performance curve: models performed best when relevant information is at the beginning or the end of the input context, but their performance declines sharply when the information is located in the middle. In this experiment, we followed procedures and data sets used in (Liu et al., 2024), to examine how token predictions are distributed as a function of input context length and position of relevant information.
In which layer(s) does the model’s correct prediction happen and finalize? As illustrated in Figure 2, the top-1 token that matches the correct answer typically emerges in the middle or later layers. When handling long input contexts with relevant information placed in the middle, the correct top-1 tokens are identified similar to earlier, but it takes many additional layers before the model finalizes the prediction. The logit lens reveals how predictions are iteratively refined as they pass through each successive layer. Early layers may produce outputs that seem plausible but are far from accurate. The model begins with rough guesses, which are gradually improved as it incorporates more context and relevant information. This study indicates that when identifying bottlenecks or inefficiencies in the model, it’s crucial to consider the position of relevant information. This positioning impacts to some extent the refinement process of token prediction, highlighting which layers on average are critical for determining the final output.
Figure 3: High-level overview of probing (Reproduced from Figure 2 in (Levinstein and Herrmann, 2023)).
4 Relation of Logit Lens to Probing
Probing refers to training a classifier to predict properties from internal representation to identify such properties, which is presumably encoded in learned representations (Alain, 2016; Ettinger et al., 2016; Hupkes et al., 2018) (see Figure 3). High probing accuracy in a layer indicates that the correct answer can be extracted from its hidden states (see Section 4.1 in (Levinstein and Herrmann, 2023)). However, this standard is often too easy to meet, particularly in straightforward classification tasks with high-dimensional hidden states (Hewitt and Liang, 2019). In contrast, a high logit lens performance signifies that the layer effectively encodes correct answers along a direction in the residual stream, providing significantly more insight as shown in interventional experiments (Li et al., 2022) that internal representations can be used to control the output of the network. Conversely, low logit lens performance does not necessarily mean that the correct answers cannot be decoded from that layer.
5 Relevance to AI Safety
Understanding inner working mechanisms of LLMs is essential for ensuring their safe development as we move toward more powerful models. Mechanistic interpretability approach has the potential to significantly advance LLM/AI safety research by providing a richer, stronger foundation for model evaluation (Casper, 2023). It can also offer early warnings of emergent capabilities, enabling us to understand better how internal structures and representations incrementally evolve as models learn, and to detect new skills or behaviors in models before they fully develop (Wei et al., 2022; Steinhardt, 2023; Nanda et al., 2023; Barak et al., 2022). Furthermore, interpretability can strengthen theoretical risk models with concrete evidence, such as identifying inner misalignment (when a model’s behavior diverges from its intended goals). By exposing potential risks or problematic behaviors, interpretability could also prompt a shift within the AI community toward adopting more rigorous safety protocols (Hubinger, 2019).
When it comes to specific AI risks (Hendrycks et al., 2023), interpretability is a powerful tool for preventing malicious misuse by identifying and eliminating sensitive information embedded in models (Meng et al., 2022; Nguyen et al., 2022). It can also alleviate competitive pressures by providing clear evidence of potential threats, fostering a culture of safety within organizations, and reinforcing AI alignment—ensuring that AI systems stay aligned with intended goals—through enhanced monitoring and evaluation (Hendrycks and Mazeika, 2022). In addition, interpretability offers essential safety checks throughout the AI development process. For example, prior to training, it informs deliberate design choices to enhance safety (Hubinger, 2019); during training, it detects early signs of misalignment, enabling proactive shifts toward alignment (Hubinger, 2022; Sharkey, 2022); and after training, it ensures rigorous evaluation of artificial cognition, verifying the model’s honesty (Burns et al., 2023; Zou et al., 2023) and screening for deceptive behaviors (Park et al., 2023). This comprehensive approach helps safeguard AI systems at every stage, driving safer and more trustworthy advancements in AI.
The emergence of internal world models in LLMs holds transformative potential for AI alignment research. If we can identify internal representations of human values and align the AI system’s objectives accordingly, achieving alignment may become a straightforward task (Wentworth, 2022). This is especially promising if the AI’s internal world model remains distinct from any notions of goals or agency (Ruthenis, 2022). In such cases, simply interpreting the world model could be enough to ensure alignment (Ruthenis, 2023), providing an effective path to safer AI systems.
Mechanistic interpretability plays a vital role in advancing various AI alignment initiatives, including the understanding and control of existing models, facilitating AI systems in solving alignment challenges, and developing robust alignment theories (technicalities and Stag, 2023; Hubinger, 2020). By enhancing strategies to detect deceptive alignment—a scenario where a model appears aligned while pursuing misaligned goals without raising suspicion (Park et al., 2023)—and eliciting latent knowledge from models (Christiano et al., 2021), we can significantly improve scalable oversight, such as through iterative distillation and amplification (Chan, 2023). Moreover, comprehensive interpretability can serve as an alignment strategy by helping us identify internal representations of human values and guiding the model to pursue those values through re-targeting an internal search process (Wentworth, 2022). Ultimately, the connection between understanding and control is crucial; a deeper understanding enables more reliable control over AI systems.
6 Discussion and Limitations
The effectiveness of the logit lens can vary significantly depending on the task and model used. For instance, tasks that require deep, abstract reasoning, layers far from the output may offer little relevant information when analyzed solely through token prediction.
Explaining how each layer of a model contributes to the final prediction by reducing complex representations to token predictions introduces biases. Many layers may serve other functions (e.g., refining internal features or re-weighting contextual information) which are not directly captured by logits. For instance, middle layers may not just be involved in predicting the next token, so using the logit lens alone may overlook their other roles in the model.
The logit lens relies on predicting which tokens have high probabilities at different layers. However, token distributions can be highly skewed, with some tokens dominating; even though they may not reflect the true nature of the underlying hidden state. This can obscure the true purpose or significance of intermediate representations.
Useful perspective on how language models construct their predictions can be obtained using logit lens. Its capabilities, however, are constrained when it comes to unveiling the model with more complex internal operations. Logit lens do not fully reveal the layered, nuanced, and abstract processes that underlie a language model’s true capabilities. Thus it is essential to utilize logit lens in combination with other techniques that focus on different aspects of model behavior (e.g., neuron activity and attention patterns), to uncover the internal workings, and to obtain a deeper and more comprehensive understanding of how language models function.
7 Conclusion
Investigating the intermediate representations by examining how a model iteratively refines token predictions, offers insights into how a model’s internal representations evolve. In this study, we utilized the logit lens to meticulously examine next token prediction refinements across different conditions, to some degree, to uncover critical nuances in how predictions are progressively refined. A deeper grasp of the prediction refinement process will enhance the development in AI safety research and development. That is, by establishing connections between the previous findings and existing hypotheses, we aim to advance the understanding of LLM behavior. These insights are vital for the development of AI technologies that are not only highly effective but also secure and aligned with human values.
Acknowledgments and Disclosure of Funding
The author would like to thank the support and opportunity given by Mukaya (Tai) Panich and Thanapong Boontaeng to research this topic. The author would also like to thank the members of the SCB 10X team for their useful comments during the contract term from May to October 2024.